Jack successfully makes a freethrow in basketball with probability 0.60. If he takes 15 freethrows, what is the variance? Round your answer to the nearest tenth.
Answer #1

Solution:-

Given that p = 0.6, n = 15

q = 1-p = 1-0.6 = 0.4

The variance = n*p*q = 15*0.6*0.4 = 3.6